# Can Internal Software Metrics Predict App Popularity at Launch? Yeas! and Nays!

### v1

@@ 3.3. Feature Extraction
System-Level Metrics. Architectural erosion—stemming from poor modularity and unmanaged dependencies—can significantly increase maintenance overhead and degrade software quality over time. To systematically quantify architectural modularity and assess design-time maintainability risks, we employ the DV8 tool (Cai and Kazman, 2019), which encapsulates decades of research in software design analysis and education (Cai and Kazman, 2023).
@@
Table 1. Seventeen System-Level Metrics Extracted Using the DV8 Tool.
@@
Table 2. Class-Level Metrics and Descriptions.
@@
String Literals Quantity Number of string literals.
Numbers Quantity Number of numeric literals.
@@
Table 3. Method-Level Metrics and Descriptions.
@@
For normal classes, we compute the minimum, maximum, mean, and 11 percentiles for all class-level features. In addition to these statistics, some metrics carry global significance. We identify seven such important global metrics: uniqueWordsQty, lambdasQty, innerClassesQty, anonymousClassesQty, loopQty, LOC, and totalMethodsQty. For these metrics, we also calculate their total count across the entire codebase for all types of classes to capture their overall impact. To complement the aggregated features, we introduce a new metric, TotalNormalClasses, which captures the number of standard classes in the app. Applications with fewer than five normal classes are excluded from the analysis to ensure statistical reliability.
@@
Table 4. Code Smells and Descriptions.
@@
Wide Hierarchy Very broad inheritance trees that may signal overly generic or complex designs.
Cyclic Dependency Cycles among classes or packages, indicating tight coupling.
@@
Unstable Dependency Modules depend on components that change frequently, reducing stability.
Hub-like Modularization A component with excessive incoming and outgoing dependencies.
Dense Structure Excessive interconnections between elements in a system, hindering modularity.
Abstract Function Call From Constructor Invoking an abstract method in a constructor, which may lead to unexpected behavior.
Multifaceted Abstraction An abstraction that combines unrelated concerns, violating the separation of concerns principle.
Scattered Functionality Related functionalities distributed across multiple locations, reducing cohesion.
Ambiguous Interface Interfaces that declare methods with unclear or conflicting responsibilities.
Cyclic Hierarchy Inheritance relationships that form a cycle, which is conceptually invalid.
Deep Hierarchy Deep inheritance chains that hinder code understanding and increase complexity.
There are many tools available for code smell detection, some of which take the APK as input (Hecht et al., 2015; SonarQube, [n. d.]), while others operate solely on source code (Palomba et al., 2017a; Moha et al., 2010; Sharma et al., 2016). For our analysis, we chose the DesigniteJava tool, which is a version of the original Designite tool (Sharma et al., 2016) specifically designed for Java. The primary reason for selecting this tool is its comprehensive support for various types of code smells, including architectural, design, implementation, and test smells. For this study, we focused on architectural, design, and implementation smells, which collectively account for a total of 34 code smells.
@@
Table 5. Metadata Description of Selected App Features.
@@
3.4. Feature Selection
Following the feature extraction process and the removal of applications for which extraction was unsuccessful, we constructed a dataset consisting of 446 Android applications, each described by 1,167 features. While this comprehensive feature set was designed to capture a wide range of structural and semantic properties of software systems, the inclusion of high-dimensional feature spaces can adversely affect the performance of machine learning models—a phenomenon commonly attributed to the curse of dimensionality. To mitigate this issue and enhance model generalizability, we applied feature selection techniques to reduce the dimensionality of the dataset while preserving relevant information. The specific subset of features employed in each experimental setting, along with their respective contributions to model performance, is discussed in detail in Section 4.
Feature Set 1: Size-only. Although numerous software metrics have been proposed and empirically evaluated as predictors of maintainability and other quality attributes, the reliability of these metrics—beyond basic size—remains a topic of ongoing debate (Chowdhury et al., 2022; Gil and Lalouche, 2017; Shepperd, 1988; Gil and Lalouche, 2016). Several studies argue that code size is the most robust and consistent predictor of maintenance effort (El Emam et al., 2001; Gil and Lalouche, 2017), while many other code metrics exhibit strong collinearity with size (Herraiz et al., 2007). Based on this evidence, we constructed a minimal baseline model using only a size-related feature to assess its standalone predictive power. To measure size, we used the CLOC (Danial, 2021) tool, which analyzes a project directory and reports the number of blank lines, comment lines, and lines of code. For each app, we computed the lines of code as size, excluding blank lines and comments.
@@ 3.4. Feature Selection
3.5. Learning Algorithms and Evaluation Metrics
@@
4. APPROACH, ANALYSIS, AND RESULTS
@@
4.1. RQ1: Can we predict an app’s rating using internal software metrics?
@@
4.1.1. Regression: Predicting the Actual Ratings
@@
Building upon these three feature sets, we conducted three groups of experiments in which six regression models were applied to predict the continuous application ratings: MLPRegressor (MLP), Lasso, Ridge, DecisionTreeRegressor (DT), RandomForestRegressor (RF), and GradientBoostingRegressor (GB). To ensure that the predicted ratings remained within the valid interval [1,5], we applied a post-processing step wherein predictions falling below 1 were clamped to 1, and those exceeding 5 were truncated to 5. This normalization was necessary due to the unbounded nature of regression outputs, which can produce values outside the expected domain. The comparative results of the regression models are summarized in Table 6.
@@
Table 6. Regression metrics (RMSE, MAE, $R^{2}$) of rating regressors evaluated using three feature selection strategies: Size-only, Handpicked, and Voting. Bold values indicate the best metric scores within each feature set.
@@
A closer examination of the regression outcomes reveals that, for the majority of cases, the RMSE remains below 1 and the MAE is under 0.6. These values suggest that the internal code metrics possess reasonable predictive capability: for instance, the Ridge regression model achieves an MAE of 0.51, indicating that it can predict application ratings with around half-point deviation on average. Such results imply that, in absolute terms, the models are able to approximate user ratings with relatively low error.
However, this positive impression is tempered by a contrasting trend in the $R^{2}$ scores. While RMSE and MAE reflect the magnitude of prediction errors, the $R^{2}$ metric evaluates the proportion of variance in the target variable that is explained by the model. While $R^{2}$ values close to 1 indicate strong explanatory power, the metric can range from $\infty$- ∞ to 1. A value of 0 means the model performs no better than predicting the mean of the target variable, while negative values indicate that the model performs worse than this baseline. In our experiments, most models yield $R^{2}$ scores that are near zero or even negative, suggesting limited ability to capture the variance in user ratings. Notably, the Ridge model—despite its low error metrics—achieves only a modest $R^{2}$ of 0.05, and several models (e.g., MLP) produce significantly negative values.
To better understand the limited performance, we analyzed the distribution of both actual and predicted ratings using the Cumulative Distribution Function (CDF), as shown in Figure 3. The CDF of the true application ratings (Figure 3(a)) reveals a significant skew in the dataset: approximately 10% of applications received rating 3, while over 5% achieved the maximum rating (5). Most notably, nearly 70% of the applications fall within the 3.0 to 4.5 range, indicating a concentrated distribution around moderately high ratings. This skew is reflected in the predicted values generated by the Ridge regression model (Figure 3(b)), where more than 95% of predictions also lie within the 3.0 to 4.5 interval.
@@
cdf of actual ratings
(a) CDF of actual ratings.
[…]
cdf of predicted ratings
(b) CDF of predicted ratings of Ridge regression model.
Figure 3. Cumulative Distribution Functions (CDFs) of actual and predicted ratings. Approximately 70% of the actual ratings fall between 3.0 and 4.5, while over 95% of the predicted ratings lie within this range, indicating the model’s bias towards predicting around the mean value.
[…]
CDF of absolute errors.
(a) CDF of absolute errors.
[…]
CDF of Squared Errors
(b) CDF of squared errors.
Figure 4. Cumulative Distribution Functions (CDFs) of absolute error and squared error for the Ridge regression model in predicting actual ratings. The plots indicate that a small number of data points account for the majority of the large errors.
[…]
Scatter Plot of True Ratings and Predicted Ratings
Figure 5. Scatter plot of true ratings and predicted ratings for each app.
Figure 6. Scatter Plot of True Ratings vs Predicted Ratings.
[…]
Scatter plot of true ratings vs predicted ratings.
In summary, the strong skew in the dataset leads the regression models to consistently predict values near the mean, resulting in low RMSE and MAE but offering limited insight into the underlying patterns of the data. This behavior is reflected in the uniformly low—often negative—$R^{2}$ values, indicating that the models fail to meaningfully capture the variance in user ratings. Consequently, the regression results do not offer conclusive evidence regarding the predictive utility of internal code metrics; they neither confirm their effectiveness nor definitively refute their relevance.
@@
4.1.2. Classification: Predicting Popularity Based on Ratings.
@@
Similar to the previous experiments, we also have three distinct feature sets for the classification task: Size-only, Handpicked, and Voting. The Voting method highlighted five key features: containsAds, genreId_TRAVEL_AND_LOCAL, parametersQty_20, innerClassesQty_50_class, and parametersQty_99. These feature sets were used to train and evaluate various classification models using Leave-One-Out Cross-Validation (LOOCV), as outlined in Section 3.5.
To address the inherent class imbalance in our dataset, we implemented three widely used resampling techniques: Random Over Sampling (ROS), Random Under Sampling (RUS), and Synthetic Minority Oversampling Technique (SMOTE) (Chawla et al., 2002). These techniques were applied to ensure that the classifier had a more balanced representation of different classes, ultimately improving model generalization. Given that we employed Leave-One-Out Cross-Validation (LOOCV) as our validation strategy, resampling was performed within each fold to prevent data leakage and maintain the integrity of the training process. This ensured that the model was evaluated on unseen data without exposure to artificially generated or resampled instances from the test fold. Additionally, the numerical features were normalized by applying standard scaling using scikit-learn (Pedregosa et al., 2011), which standardizes each feature by removing its mean and scaling it to unit variance. This normalization is particularly beneficial for models that rely on gradient-based optimization, such as neural networks.
The results of our classification experiments, presented in Table 7, mirror trends observed in the regression setting: models trained on the Voting feature set consistently outperform those utilizing either the Size-only or Handpicked feature sets. In particular, models trained solely on the Size-only features exhibit the weakest performance, with Matthews Correlation Coefficient (MCC) values that are close to or below zero. This indicates poor discriminative ability, often comparable to or worse than random guessing.
Table 7. Performance metrics (Macro F1, MCC, AUC) of rating-based popularity classifiers evaluated using three feature selection strategies: Size-only, Handpicked, and Voting. Bold values indicate the best metric scores within each feature set.
@@
Overall, the MLPClassifier trained with SMOTE on the Voting feature set achieved the strongest and most balanced performance across all evaluated metrics. The model attained a Macro F1 Score of 0.72, indicating a balanced classification performance across both classes. Furthermore, the Matthews Correlation Coefficient (MCC) improved to 0.45, demonstrating a meaningful increase in overall predictive reliability. The Area Under the Receiver Operating Characteristic Curve (AUC-ROC) also reached 0.72, suggesting that the model was more capable of distinguishing between the positive and negative classes.
The class-wise performance of the SMOTE-enhanced MLPClassifier, shown in Table 8, highlights its effectiveness in mitigating class imbalance. The model achieved a precision of 0.567, a recall of 0.639, and f1-score of 0.601 for the Unpopular class, indicating improved minority class recognition while still facing some misclassification challenges. Meanwhile, the Popular class demonstrated stronger performance with a precision of 0.862, a recall of 0.823, and f1-score of 0.842, ensuring high reliability in identifying high-rated apps. These results confirm that SMOTE effectively improved recall for the minority class while maintaining high precision for the majority class, leading to a Macro F1-Score of 0.72 and MCC of 0.45. By generating synthetic examples, SMOTE allowed the model to learn better decision boundaries, enhancing overall classification performance. The confusion matrix and AUC-ROC curve in Figure 7 further illustrate the classifier’s improved ability to separate Popular from Unpopular apps, as reflected in a higher true positive rate for the minority class and an AUC score of 0.72.
Table 8. Classification report for rating-based popularity prediction using MLP classifier with SMOTE.
@@
[…]
Confusion Matrix and AUC-ROC Curve (App Rating)
Figure 7. Confusion matrix and AUC-ROC curve for rating-based popularity prediction using MLP classifier with SMOTE.
@@
Summary of RQ1 Models trained on statistically selected (Voting) features consistently outperformed those using Size-only or Handpicked features across both regression and classification tasks. While regression models achieved low error rates, their explanatory power was limited due to the strong skew in the rating distribution, which caused predictions to cluster around the mean. In contrast, classification models, especially MLPClassifier with SMOTE, performed better at distinguishing between Popular and Unpopular apps, highlighting the predictive value of internal software metrics when appropriately framed.
4.2. RQ2: Can we predict an app’s number of downloads using internal software metrics?
4.2.1. Regression: Predicting the Actual DownloadsPerYear.

### v2

@@ 3.3 Feature Extraction
System-Level Metrics. Architectural erosion—stemming from poor modularity and unmanaged dependencies—can significantly increase maintenance overhead and degrade software quality over time. To systematically quantify architectural modularity and assess design-time maintainability risks, we employ the DV8 tool (Cai and Kazman, 2019), which encapsulates decades of research in software design analysis and education (Cai and Kazman, 2023a).
@@
Table 1: Seventeen System-Level Metrics Extracted Using the DV8 Tool.
@@
Table 2: Class-Level Metrics and Descriptions.
@@
Table 3: Method-Level Metrics and Descriptions.
@@
For normal classes, we compute the minimum, maximum, mean, and 11 percentiles for all class-level features. In addition to these statistics, some metrics carry global significance. We identify seven such important global metrics: lambdasQty, totalMethodsQty, innerClassesQty, LOC, anonymousClassesQty, loopQty, and uniqueWordsQty. For these metrics, we also calculate their total count across the entire codebase for all types of classes to capture their overall impact. To complement the aggregated features, we introduce a new metric, TotalNormalClasses, which captures the number of standard classes in the app. Applications with fewer than five normal classes are excluded (57 applications) from the analysis to ensure statistical reliability.
@@
Table 4: A Subset of Code Smells and Descriptions.
@@
There are many tools available for code smell detection, some of which take the APK as input (Hecht et al., 2015; SonarQube, 2023), while others operate solely on source code (Palomba et al., 2017a; Moha et al., 2010; Sharma et al., 2016). For our analysis, we chose the DesigniteJava tool, which is a version of the original Designite tool (Sharma et al., 2016) specifically designed for Java. The primary reason for selecting this tool is its comprehensive support for various types of code smells, including architectural, design, implementation, and test smells. For this study, we focused on architectural, design, and implementation smells, which collectively account for a total of 34 code smells.
@@
Table 5: Metadata Description of Selected App Features.
@@
3.4 Feature Selection
After excluding 57 applications from the initial pool of 503 during class-level code metrics extraction, we obtained a final dataset of 446 Android applications, each characterized by 1,167 features. While this comprehensive feature set was designed to capture a wide range of structural and semantic properties of software systems, the inclusion of high-dimensional feature spaces can adversely affect the performance of machine learning models—a phenomenon commonly attributed to the curse of dimensionality. To mitigate this issue and enhance model generalizability, we applied feature selection techniques to reduce the dimensionality of the dataset while preserving relevant information. The specific subset of features employed in each experimental setting, along with their respective contributions to model performance, is discussed in detail in Section 4.
Feature Set 1: Size-only. Although numerous software metrics have been proposed and empirically evaluated as predictors of maintainability and other quality attributes, the reliability of these metrics, beyond basic size, remains a topic of ongoing debate (Chowdhury et al., 2022; Gil and Lalouche, 2017; Shepperd, 1988; Gil and Lalouche, 2016). Several studies argue that the size of code is the most robust and consistent predictor of maintenance effort (El Emam et al., 2001; Gil and Lalouche, 2017), while many other code metrics exhibit strong collinearity with size (Herraiz et al., 2007). Based on this evidence, we constructed a minimal baseline model using only a size-related feature to assess its standalone predictive power. To measure size, we used the CLOC (Danial, 2021) tool, which analyzes a project directory and reports the number of blank lines, comment lines, and lines of code. For each app, we computed the lines of code as size, excluding blank lines and comments.
@@ 3.4 Feature Selection
3.5 Learning Algorithms and Evaluation Metrics
@@
4 APPROACH, ANALYSIS, AND RESULTS
@@
4.1 RQ1: To what extent can internal software metrics predict an app’s user rating?
@@
4.1.1 Regression: Predicting the Actual Ratings
@@
Building upon these three feature sets, we conducted three groups of experiments in which six regression models were applied to predict the continuous app ratings: MLPRegressor (MLP), Lasso, Ridge, DecisionTreeRegressor (DT), RandomForestRegressor (RF), and GradientBoostingRegressor (GB). To ensure that the predicted ratings remained within the valid interval [1,5], we applied a post-processing step wherein predictions falling below 1 were clamped to 1, and those exceeding 5 were truncated to 5. This normalization was necessary due to the unbounded nature of regression outputs, which can produce values outside the expected domain. The comparative results of the regression models are summarized in Table 6.
@@
Table 6: Regression metrics (RMSE, MAE, $R^{2}$) of rating regressors evaluated using three feature selection strategies: Size-only, Handpicked, and Voting. Bold values indicate the best metric scores within each feature set.
@@
A closer examination of the regression outcomes reveals that, for the majority of cases, the RMSE remains below 1 and the MAE is under 0.6. These values may suggest that the internal code metrics possess reasonable predictive capability: for instance, the Ridge regression model achieves an MAE of 0.51, indicating that it can predict application ratings with around half-point deviation on average. Such results imply that, in absolute terms, the models are able to approximate user ratings with relatively low error.
However, this positive impression is tempered by a contrasting trend in the $R^{2}$ scores. While RMSE and MAE reflect the magnitude of prediction errors, the $R^{2}$ metric evaluates the proportion of variance in the target variable that is explained by the model. While $R^{2}$ values close to 1 indicate strong explanatory power, the metric can range from -$\infty$ to 1. A value of 0 means the model performs no better than predicting the mean of the target variable, while negative values indicate that the model performs worse than this baseline. In our experiments, most models yield $R^{2}$ scores that are near zero or even negative, suggesting limited ability to capture the variance in user ratings. Notably, the Ridge model—despite its low error metrics—achieves only a modest $R^{2}$ of 0.05, and several models (e.g., MLP) produce significantly negative values.
(a) CDF of actual ratings.
(b) CDF of predicted ratings of Ridge regression model.
Figure 3: Cumulative Distribution Functions (CDFs) of actual and predicted ratings. Approximately 70% of the actual ratings fall between 3.0 and 4.5, while over 95% of the predicted ratings lie within this range, indicating the model’s bias towards predicting around the mean value.
To better understand the limited performance, we analyzed the distribution of both actual and predicted ratings using the Cumulative Distribution Function (CDF), as shown in Figure 3. The CDF of the true application ratings (Figure 3a) reveals a significant skew in the dataset: approximately 10% of applications received rating 3, while over 5% achieved the maximum rating (5). Most notably, nearly 70% of the applications fall within the 3.0 to 4.5 range, indicating a concentrated distribution around moderately high ratings. This skew is reflected in the predicted values generated by the Ridge regression model (Figure 3b), where more than 95% of predictions also lie within the 3.0 to 4.5 interval.
(a) CDF of absolute errors.
(b) CDF of squared errors.
Figure 4: Cumulative Distribution Functions (CDFs) of absolute error and squared error for the Ridge regression model in predicting actual ratings. The plots indicate that a small number of data points account for the majority of the large errors.
Figure 5: Scatter plot of true ratings and predicted ratings for each app.
Figure 6: Scatter Plot of True Ratings vs Predicted Ratings.
@@
4.1.2 Classification: Predicting Popularity Based on Ratings.
@@
Similar to the previous experiments, we also have three distinct feature sets for the classification task: Size-only, Handpicked, and Voting. The Voting method highlighted five key features: containsAds, innerClassesQty_50_class, genreId_TRAVEL_AND_LOCAL, parametersQty_20, and parametersQty_99. These feature sets were used to train and evaluate various classification models using Leave-One-Out Cross-Validation (LOOCV), as outlined in Section 3.5.
To address the inherent class imbalance in our dataset, we implemented three widely used resampling techniques: Random Over Sampling (ROS), Random Under Sampling (RUS), and Synthetic Minority Oversampling Technique (SMOTE) (Chawla et al., 2002). These techniques were applied to ensure that the classifier had a more balanced representation of different classes, ultimately improving model generalization. Given that we employed LOOCV as our validation strategy, resampling was performed within each fold to prevent data leakage and maintain the integrity of the training process. This ensured that the model was evaluated on unseen data without exposure to artificially generated or resampled instances from the test fold. Additionally, the numerical features were normalized by applying standard scaling using scikit-learn (Pedregosa et al., 2011), which standardizes each feature by removing its mean and scaling it to unit variance. This normalization is particularly beneficial for models that rely on gradient-based optimization, such as neural networks.
The results of our classification experiments, presented in Table 7, mirror trends observed in the regression setting: models trained on the Voting feature set consistently outperform those utilizing either the Size-only or Handpicked feature sets. In particular, models trained solely on the Size-only features exhibit the weakest performance, with MCC values that are close to or below zero. This indicates poor discriminative ability, often comparable to or worse than random guessing.
Table 7: Performance metrics (Macro F1, MCC, AUC) of rating-based popularity classifiers evaluated using three feature selection strategies: Size-only, Handpicked, and Voting. Bold values indicate the best metric scores within each feature set.
@@
Overall, the MLPClassifier trained with SMOTE on the Voting feature set achieved the strongest and most balanced performance across all evaluated metrics. The model attained a Macro F1 score of 0.72, indicating a balanced classification performance across both classes. Furthermore, the MCC increased to 0.45, indicating a meaningful improvement in overall predictive reliability and suggesting that the comparatively high F1 score was not due to chance. The Area Under the Receiver Operating Characteristic Curve (AUC-ROC) also reached 0.72, suggesting that the model was more capable of distinguishing between the positive and negative classes.
The class-wise performance of the SMOTE-enhanced MLPClassifier, shown in Table 8, highlights its effectiveness in mitigating class imbalance. The model achieved an F1 score of 0.601 for the Unpopular class, indicating improved minority class recognition while still facing some misclassification challenges. Meanwhile, the Popular class demonstrated stronger performance with a precision of 0.862, a recall of 0.823, and an F1 score of 0.842, ensuring high reliability in identifying high-rated apps. These results confirm that SMOTE effectively improved recall for the minority class while maintaining high precision for the majority class, leading to a Macro F1 score of 0.72 and MCC of 0.45. By generating synthetic examples, SMOTE allowed the model to learn better decision boundaries, enhancing overall classification performance. The confusion matrix and AUC-ROC curve in Figure 7 further illustrate the classifier’s improved ability to separate Popular from Unpopular apps, as reflected in a higher true positive rate for the minority class and an AUC score of 0.72.
Table 8: Classification performance for rating-based popularity prediction using MLP classifier with SMOTE.
@@
Figure 7: Confusion matrix and AUC-ROC curve for rating-based popularity prediction using MLP classifier with SMOTE.
@@
Summary of RQ1 Due to the skewed data, the regression models did not learn much and made predictions around the mean only. As a result, we were unable to draw meaningful conclusions about the usefulness of internal metrics from the regression results alone. In contrast, classification models—especially MLPClassifier with SMOTE—demonstrated significantly better performance in distinguishing between Popular and Unpopular apps, highlighting the predictive value of internal software metrics when the problem is appropriately framed.
4.2 RQ2: To what extent can internal software metrics predict an app’s number of downloads?
4.2.1 Regression: Predicting the Actual DownloadsPerYear.